### v3

@@ Appendix B Missing Proofs in Section 4
where $g^{t}\coloneqq\mathbb{\mathbb{E}}\left[\widehat{g}^{t}|\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$ and $\xi^{t}=\widehat{g}^{t}-g^{t}$. Next, we bound these four terms respectively.
@@ IV $\leq LD_{\psi}(x^{t+1},x^{t})+M\sqrt{2D_{\psi}(x^{t+1},x^{t})}$
$\gamma_{T}v_{T}\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\right])+\sum_{t=1}^{T}\gamma_{t}v_{t-1}\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]$
@@
where the last line is due to $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\right]=\mathbb{\mathbb{E}}\left[\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{2}\mid\mathcal{F}^{t-1}\right]\right]\leq\sigma^{2}$ (Assumption 5A.) and $\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\left\langle\mathbb{\mathbb{E}}\left[\xi^{t}|\mathcal{F}^{t-1}\right],z^{t-1}-x^{t}\right\rangle\right]=0$ ($z^{t-1}-x^{t}\in\mathcal{F}^{t-1}$ and Assumption 4.). Finally, we divide both sides by $\gamma_{T}v_{T}$ and plug in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}=\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}$ to finish the proof. ∎
@@
$\mathbb{\mathbb{E}}\left[U_{t}\mid\mathcal{F}^{t-1}\right]$ $=U_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(2w_{t}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}-2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\right)\mid\mathcal{F}^{t-1}\right]$
@@
Hence, we know $\mathbb{\mathbb{E}}\left[U_{T}\right]\leq U_{0}=1$. Thus, there is
$\Pr\left[U_{T}>\frac{2}{\delta}\right]\overset{(b)}{\leq}\frac{\delta}{2}\mathbb{\mathbb{E}}\left[U_{T}\right]\leq\frac{\delta}{2}$
@@
$\mathbb{\mathbb{E}}\left[R_{t}\mid\mathcal{F}^{t-1}\right]$ $=R_{t-1}\mathbb{\mathbb{E}}\left[\exp\left(w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle-w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}\right)\mid\mathcal{F}^{t-1}\right]$
@@
where $(c)$ is by applying Lemma 2.1 (note that $z^{t-1}-x^{t}\in\mathcal{F}^{t-1}=\sigma(\widehat{g}^{s},s\in\left[t-1\right])$). Hence, we have $\mathbb{\mathbb{E}}\left[R_{T}\right]\leq R_{0}=1$, which immediately implies
$\Pr\left[R_{T}>\frac{2}{\delta}\right]\overset{(d)}{\leq}\frac{\delta}{2}\mathbb{\mathbb{E}}\left[R_{T}\right]\leq\frac{\delta}{2}$
@@ Appendix C General Convex Functions
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})}\log T}{\sqrt{T}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})\log T}}{\sqrt{T}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (19)
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (20)
@@
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\eta\sum_{t=1}^{T}1/\sqrt{t}}+2\eta(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t\sum_{s=t}^{T}1/\sqrt{s}}$
@@
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T/2L}+\frac{M^{2}+\sigma^{2}}{L}\sum_{t=1}^{T}\frac{1}{T-t+1}=\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{M^{2}+\sigma^{2}}{L}\sum_{t=1}^{T}\frac{1}{t}$
@@
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}$
@@
    $\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T^{2}}\left(2LT+\frac{5T^{\frac{3}{2}}}{3\eta}\right)+2(M^{2}+\sigma^{2})\left[\frac{2\eta(1+\log T)}{\sqrt{T}}+\frac{2\eta(1+\log T)}{\sqrt{T}}\right]$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+4\eta(M^{2}+\sigma^{2})(1+\log T)\right]\lor\left[\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2\eta(M^{2}+\sigma^{2})(1+\log T)}{\sqrt{T}}\right]$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T}\left(2L\lor\frac{\sqrt{T}}{\eta}\right)+2\left(\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}\right)(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{T-t+1}$
@@
C.1 Optimal Rates via the Linear Decay Step Size
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\right]\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})}}{\sqrt{T}}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (31)
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (32)
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}T-t+1}\left(2LT\lor\frac{T^{\frac{3}{2}}}{\eta}\right)+\frac{2\eta(M^{2}+\sigma^{2})}{T^{\frac{3}{2}}}\sum_{t=1}^{T}\frac{(T-t+1)^{2}}{\sum_{s=t}^{T}T-s+1}$
@@ D.1 The Case of $\mu_{f}>0$
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\begin{cases}O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{L(1+\kappa_{f})D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t=1\\ \frac{1}{\mu_{f}(\eta+2\kappa_{f})}&2\leq t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ where $\eta\geq 0$ can be any number satisfying $\eta+\kappa_{f}>1$ and $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2\eta+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
@@
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}.$ (36)
Now observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\eta_{t}\Gamma_{t}=\begin{cases}\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{f}}&t\geq 2\\ \eta_{1}&t=1\end{cases},$
where $\Gamma_{t\in\left[T\right]}\coloneqq\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}$. Hence, (36) can be rewritten as
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\eta_{1}^{2}}{\eta_{1}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\eta_{t}^{2}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\eta_{t})}\right].$ (37)
Now let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ for our three choices respectively:
@@
First, we consider $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$. We can calculate $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\Gamma_{t\in\left[T\right]}=\prod_{s=2}^{t}\frac{1}{1-\mu_{f}\eta_{s}}=\prod_{s=2}^{t}\frac{s+2\kappa_{f}}{s-1+2\kappa_{f}}=\frac{t+2\kappa_{f}}{1+2\kappa_{f}}.$
Hence, using (37), we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{(1-\frac{1}{1+2\kappa_{f}})D_{\psi}(x,x^{1})}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+2(M^{2}+\sigma^{2})\left[\frac{\frac{1}{\mu_{f}^{2}(1+2\kappa_{f})^{2}}}{\frac{1}{\mu_{f}(1+2\kappa_{f})}+\frac{\Gamma_{T}-1}{\mu_{f}}}+\sum_{t=2}^{T}\frac{\mu_{f}\cdot\frac{1}{\mu_{f}^{2}(t+2\kappa_{f})^{2}}}{(\Gamma_{T}/\Gamma_{t-1}-1)(1-\mu_{f}\cdot\frac{1}{\mu_{f}(t+2\kappa_{f})})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{(1+2\kappa_{f})\Gamma_{T}-2\kappa_{f}}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{1}{(1+2\kappa_{f})((1+2\kappa_{f})\Gamma_{T}-2\kappa_{f})}+\sum_{t=2}^{T}\frac{1}{(\Gamma_{T}/\Gamma_{t-1}-1)(t-1+2\kappa_{f})(t+2\kappa_{f})}\right]$
$=$ $\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2(M^{2}+\sigma^{2})}{\mu_{f}}\left[\frac{1}{(1+2\kappa_{f})T}+\sum_{t=2}^{T}\frac{1}{(T-t+1)(t+2\kappa_{f})}\right]$

### v4

@@ Appendix B Missing Proofs in Section 4
where $g^{t}\coloneqq\mathbb{E}\left[\widehat{g}^{t}|\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$ and $\xi^{t}=\widehat{g}^{t}-g^{t}$. Next, we bound these four terms respectively.
@@ IV $\leq LD_{\psi}(x^{t+1},x^{t})+M\sqrt{2D_{\psi}(x^{t+1},x^{t})}$
$\gamma_{T}v_{T}\mathbb{E}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $(1-\mu_{f}\eta_{1})v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}2\gamma_{t}\eta_{t}v_{t}(M^{2}+\mathbb{E}\left[\left\|\xi^{t}\right\|_{*}^{2}\right])+\sum_{t=1}^{T}\gamma_{t}v_{t-1}\mathbb{E}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]$
@@
where the last line is due to $\mathbb{E}\left[\left\|\xi^{t}\right\|_{*}^{2}\right]\leq\sigma^{2}$ (Assumption 5A.) and $\mathbb{E}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]=\mathbb{E}\left[\left\langle\mathbb{E}\left[\xi^{t}|\mathcal{F}^{t-1}\right],z^{t-1}-x^{t}\right\rangle\right]=0$ ($z^{t-1}-x^{t}\in\mathcal{F}^{t-1}$ and Assumption 4.). Finally, we divide both sides by $\gamma_{T}v_{T}$ and plug in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{w_{T}\gamma_{T}}{\sum_{s=t\lor 1}^{T}w_{s}\gamma_{s}}=\frac{\gamma_{T}}{\sum_{s=t\lor 1}^{T}\gamma_{s}}$ to finish the proof. ∎
@@
$\mathbb{E}\left[U_{t}\mid\mathcal{F}^{t-1}\right]$ $=U_{t-1}\mathbb{E}\left[\exp\left(2w_{t}\gamma_{t}\eta_{t}v_{t}\left\|\xi^{t}\right\|_{*}^{2}-2w_{t}\gamma_{t}\eta_{t}v_{t}\sigma^{2}\right)\mid\mathcal{F}^{t-1}\right]$
@@
Hence, we know $\mathbb{E}\left[U_{T}\right]\leq U_{0}=1$. Thus, there is
$\Pr\left[U_{T}>\frac{2}{\delta}\right]\overset{(b)}{\leq}\frac{\delta}{2}\mathbb{E}\left[U_{T}\right]\leq\frac{\delta}{2}$
@@
$\mathbb{E}\left[R_{t}\mid\mathcal{F}^{t-1}\right]$ $=R_{t-1}\mathbb{E}\left[\exp\left(w_{t}\gamma_{t}v_{t-1}\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle-w_{t}^{2}\gamma_{t}^{2}v_{t-1}^{2}\sigma^{2}\left\|z^{t-1}-x^{t}\right\|^{2}\right)\mid\mathcal{F}^{t-1}\right]$
@@
where $(c)$ is by applying Lemma 2.1 (note that $z^{t-1}-x^{t}\in\mathcal{F}^{t-1}=\sigma(\widehat{g}^{s},s\in\left[t-1\right])$). Hence, we have $\mathbb{E}\left[R_{T}\right]\leq R_{0}=1$, which immediately implies
$\Pr\left[R_{T}>\frac{2}{\delta}\right]\overset{(d)}{\leq}\frac{\delta}{2}\mathbb{E}\left[R_{T}\right]\leq\frac{\delta}{2}$
@@ Appendix C General Convex Functions
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})}\log T}{\sqrt{T}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\log T\right]\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})\log T}}{\sqrt{T}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (19)
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (20)
@@
    $\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\eta\sum_{t=1}^{T}1/\sqrt{t}}+2\eta(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t\sum_{s=t}^{T}1/\sqrt{s}}$
@@
    $\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T/2L}+\frac{M^{2}+\sigma^{2}}{L}\sum_{t=1}^{T}\frac{1}{T-t+1}=\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{M^{2}+\sigma^{2}}{L}\sum_{t=1}^{T}\frac{1}{t}$
@@
    $\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}$
@@
    $\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T^{2}}\left(2LT+\frac{5T^{\frac{3}{2}}}{3\eta}\right)+2(M^{2}+\sigma^{2})\left[\frac{2\eta(1+\log T)}{\sqrt{T}}+\frac{2\eta(1+\log T)}{\sqrt{T}}\right]$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+4\eta(M^{2}+\sigma^{2})(1+\log T)\right]\lor\left[\frac{2LD_{\psi}(x,x^{1})}{T}+\frac{2\eta(M^{2}+\sigma^{2})(1+\log T)}{\sqrt{T}}\right]$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{T}\left(2L\lor\frac{\sqrt{T}}{\eta}\right)+2\left(\frac{1}{2L}\land\frac{\eta}{\sqrt{T}}\right)(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{1}{T-t+1}$
@@
C.1 Optimal Rates via the Linearly Decaying Step Size
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{1}{\sqrt{T}}\left[\frac{D_{\psi}(x,x^{1})}{\eta}+\eta(M^{2}+\sigma^{2})\right]\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M+\sigma)\sqrt{D_{\psi}(x,x^{1})}}{\sqrt{T}}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (31)
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$ (32)
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}T-t+1}\left(2LT\lor\frac{T^{\frac{3}{2}}}{\eta}\right)+\frac{2\eta(M^{2}+\sigma^{2})}{T^{\frac{3}{2}}}\sum_{t=1}^{T}\frac{(T-t+1)^{2}}{\sum_{s=t}^{T}T-s+1}$
@@
$F(x^{T+1})-F(x)\leq\frac{4D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+4\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=t}^{T}\eta_{s}}.$
@@ D.1 The Case of $\mu_{f}>0$
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\begin{cases}O\left(\frac{LD_{\psi}(x,x^{1})}{T}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}\\ O\left(\frac{L(1+\kappa_{f})D_{\psi}(x,x^{1})}{T(T+\kappa_{f})}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right)&\eta_{t\in\left[T\right]}=\frac{2}{\mu_{f}(t+1+4\kappa_{f})}\end{cases}.$
If $T$ is known, by taking $\eta_{t\in\left[T\right]}=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}&t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}&t\geq\tau+1\end{cases}$ with $\tau\coloneqq\left\lceil\frac{T}{2}\right\rceil$, there is
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq O\left(\frac{LD_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2+4\kappa_{f}}\right)}+\frac{(M^{2}+\sigma^{2})\log T}{\mu_{f}(T+\kappa_{f})}\right).$
@@
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{(1-\mu_{f}\eta_{1})D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\gamma_{t}}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}},$ (35)
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}$. Now, let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}$ for the three choices respectively:
@@
$\eta_{t\in\left[T\right]}$ $=\begin{cases}\frac{1}{\mu_{f}(1+2\kappa_{f})}=\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t\leq\tau\\ \frac{2}{\mu_{f}(t-\tau+2+4\kappa_{f})}\leq\frac{1}{\mu_{f}+2L}\leq\frac{1}{2L\lor\mu_{f}}&t\geq\tau+1\end{cases}.$
@@
First, we consider $\eta_{t\in\left[T\right]}=\frac{1}{\mu_{f}(t+2\kappa_{f})}$. We can find $\eta_{1}=\frac{1}{\mu_{f}(1+2\kappa_{f})}$ and
$\gamma_{t\in\left[T\right]}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu_{f}\eta_{s})}=\frac{1}{\mu_{f}(t+2\kappa_{f})\prod_{s=2}^{t}(1-\frac{1}{s+2\kappa_{f}})}=\frac{1}{\mu_{f}(1+2\kappa_{f})}=\eta_{1}.$ (36)
Hence, using (35) and (36), we have
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{(\eta_{1}^{-1}-\mu_{f})D_{\psi}(x,x^{1})}{T}+2(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}}{T-t+1}$